OpenStudy (anonymous):

x^2+9x-23 what is the vertex?

OpenStudy (anonymous):

x=-b/(2a) x=-9/2 = -4.5 (-4.5, (-4.5,-43.25)

OpenStudy (anonymous):

thanks, and what would the axis of symmetry be?

OpenStudy (anonymous):

-4.5

OpenStudy (anonymous):

yes, x=-4.5

OpenStudy (anonymous):

and it has a maximum?

OpenStudy (anonymous):

and where is the maximum located?

OpenStudy (anonymous):

like value?

OpenStudy (anonymous):

there isnt a max. min = -43.25. it is a parabola

OpenStudy (anonymous):

the y value of the vertex is either the max or min depending on whether it opens up or down. Since a is +1 it open up and -43.25 is the minimum and there is no maximum